### v1

# Crude Approximation of Directed Minimum Cut and Arborescence Packing in Almost Linear Time

Abstract
We give almost-linear-time algorithms for approximating rooted minimum cut and maximum arborescence packing in directed graphs, two problems that are dual to each other [edmonds1973edge]. More specifically, for an $n$-vertex, $m$-edge directed graph $G$ whose $s$-rooted minimum cut value is $k$,
  1. 1.
    Our first algorithm computes an $s$-rooted cut of size at most $O(k\log^{5}n)$ in $m^{1+o(1)}$ time, and
  2. 2.
    Our second algorithm packs $k$ $s$-rooted arborescences with $n^{o(1)}$ congestion in $m^{1+o(1)}$ time, certifying that the $s$-rooted minimum cut is at least $k/n^{o(1)}$.
Our first algorithm works for weighted graphs as well.
Prior to our work, the fastest algorithms for computing the $s$-rooted minimum cut were exact but required super-linear running time: either $\tilde{O}(mk)$ [gabow1991matroid] or $\tilde{O}(m^{1+o(1)}\min\{\sqrt{n},n/m^{1/3}\})$ [cen2022minimum]. The fastest known algorithms for packing $s$-rooted arborescences had no congestion but required $\tilde{O}(m\cdot\mathrm{poly}(k))$ time [bhalgat2008fast].
1 Introduction
The (global) minimum cut problem is very well-studied in graph algorithms. In the context of directed graphs, the goal is to find a cut (S,VS)(S,V-S) in the graph $G=(V,E)$ such that the number of edges from $S$ to VSV-S is minimized. Given an arbitrary source $s\in V$, the $s$-rooted minimum cut is a cut $(S,V\setminus S)$ of minimum size with $s\in S$. The size of the $s$-rooted minimum cut is called $s$-rooted connectivity. Observe that the global minimum cut problem can be solved by running two $s$-rooted minimum cut computations, one on the original graph and one on the graph with all edges reversed.
There is a line of work on solving the $s$-rooted minimum cut problem. Hao and Orlin [hao1994faster] showed that a $\tilde{O}(mn)$-time algorithm based on push-relabel methods [goldberg1988new]. Then, Gabow [gabow1991matroid] gave an algorithm with $\tilde{O}(mk)$ time where $k$ is the size of the $s$-rooted minimum cut. A recent work [cen2022minimum] solves the problem in $\min(n/m^{1/3},\sqrt{n})\cdot m^{1+o(1)}$ time.
A closely related object to $s$-rooted mincut is an $s$-rooted arborescence. An $s$-rooted arborescence is a directed tree rooted at $s$ where all arrows point away from the root $s$. A very influential result by Edmonds [edmonds1973edge] shows that the maximum number of disjoint $s$-rooted arborescences equals the size of $s$-rooted minimum cut.
In the arborescence packing problem, we need to compute the maximum number of disjoint arborescences rooted at $s$. Gabow [gabow1991matroid] an $O(n^{2}k^{2})$-time algorithm. Later, Bhalgat et al. [bhalgat2008fast] improved the running time for arborescence packing to $\tilde{O}(m\mathrm{poly}(k))$. In weighted graphs, $k$ can be very large. In this case, the fastest algorithm is by Gabow and Manu [gabow1998packing] and takes $\tilde{O}(mn^{2})$ time.
In short, the state-of-the-art algorithms for both problems still require super-linear time. It was unclear if a speed-up is possible even when we allow approximation.
1.1 Our Results
In this paper, we present the first almost-linear-time algorithms for both problems, via directed expander hierarchies.
For the $s$-rooted minimum cut problem, we give an algorithm that achieves with high probability an $O(\log^{5}(n))$-approximation.
Theorem 1.1.
Let $G=(V,E)$ be a directed graph with a given source vertex $s$ where the $s$-rooted connectivity is $k$. There is an algorithm that computes a cut $(S,V\setminus S)$ where $s\in S$ of size $O(k\cdot\log^{5}(n))$. The running time is $m^{1+o(1)}$.
Theorem 1.1 generalizes seamlessly to weighted graphs as well.
Before stating our result on arborescence packing, we first define congestion. Let ${\cal T}=\{T_{1},\dots,T_{k^{\prime}}\}$ be a set of $s$-rooted arborescences. The congestion $\cal T$} is the maximum number of times any edge may appear in $T_{i}\in{\cal T}$, i.e., $\max_{e\in E}|\{T_{i}\in{\cal T}\mid e\in T_{i}\}|$.
In this work, we design an algorithm that achieves $n^{o(1)}$ congestion in almost-linear time. [^1] We note that our proof of Theorem 1.2 (see Theorem 5.2 for a detailed version) assumes an almost-linear-time directed expander routing subroutine (see Theorem 5.1), which is commonly believed, but we are not aware of a formal proof in the existing literature. We will work on a full proof in the next version.
Theorem 1.2.
Let $G=(V,E)$ be a directed graph with a given source vertex $s$ and a parameter $k>0$. There is an algorithm that computes either
  • a cut $(S,V\setminus S)$ where $s\in S$ of size less than $k$, or
  • $k$ arborescences rooted at $s$ with congestion $n^{o(1)}$, certifying that $s$-rooted connectivity is at least $k/n^{o(1)}$.
The running time is $m^{1+o(1)}$.
Theorem 1.2 strengthens Theorem 1.1 in the senses that it also returns a cerficate of the lower bound for $s$-rooted connectivity, but the approximation factor is $n^{o(1)}$ instead of $\mathrm{polylog}(n)$.
Concurrent Work
Independently, Quanrud [quanrud2025approximatingdirectedconnectivityalmostlinear] very recently showed a $(1+\epsilon)$-approximation algorithm for $s$-rooted connectivity using $m^{1+o(1)}$ time. His approximation of $(1+\epsilon)$ is stronger than our polylogarithmic approximation factor in Theorem 1.1. It is unclear if [quanrud2025approximatingdirectedconnectivityalmostlinear] implies any result in arborescence packing similar to Theorem 1.2.
2 Preliminaries
Given a directed graph $G=(V,E)$, we denote $n=|V|$ and $m=|E|$. We use $[W]$ to abbreviate the set $\{1,\dots,W\}$. We use the notation $\tilde{O}$ to hide polylogarithmic factors in time complexity. For any set $S$ and function $f:S\to\mathbb{R}$, unless otherwise stated, for any subset $T\subseteq S$, we define $f(T)=\sum_{t\in T}f(t)$.
Directed Graphs and Connectivity.
In a directed graph $G=(V,E)$, for any set $S\subseteq V$, call (S,VS)(S,V-S) a cut. We define $E^{+}(S)$ as the edges from $S$ to $V\setminus S$, and $E^{-}(S)$ as the edges from $V\setminus S$ to $S$. We define $\delta(S)=|E^{+}(S)|$ and $\rho(S)=|E^{-}(S)|$. When $S=\{s\}$ is a singleton, we abbreviate as $\delta(s)=|E^{+}(s)|$ and $\rho(s)=|E^{-}(s)|$. The degree of $s$ is $\deg(s)=\delta(s)+\rho(s)$ and the volume of a set $S$ is $\mathrm{vol}(S)=\deg(S)=\sum_{s\in S}\deg(s)$. We can similarly discuss degree (or volume) for a subgraph $G^{\prime}$ or an edge subset $X\subseteq E$, which we denote by $\deg_{G^{\prime}}$ or $\deg_{X}$, respectively.
The (global) min-cut of a directed graph $G$ is defined as $\min_{\emptyset\neq S\subsetneq V}\rho(S)$. Given a fixed source $s\in V$, we can also define the $s$-rooted min-cut problem, which asks $\min_{\emptyset\neq S\subseteq V\setminus s}\rho(S)$. Let $G^{R}$ be the graph obtained by inverting all edges in $G$. One can observe that, the min-cut of $G$ can be computed by taking the minimum from both $s$-rooted min-cut in $G$ and $G^{R}$. Therefore, below we consider the $s$-rooted min-cut problem, and we assume $E^{-}(s)=\emptyset$ by deleting the incoming edges to $s$, which does not influence the result.
Strongly Connected Components.
For a directed graph $G$, we use $[G]$ to denote the collection of strongly connected components (SCCs) in $G$, where each SCC $C$ is represented by a subset of vertices. Therefore, $[G]$ forms a partition of $V(G)$.
Max-Flows and Min-Cuts.
Given a directed graph $G=(V,E)$ and $s,t\in V$, the max-flow from $s$ to $t$ is the maximum number of edge-disjoint paths from $s$ to $t$. By duality, this equals the size of min-cut from $s$ to $t$, which is $\min_{t\in T\subseteq V-s}\rho(T)$. Here, for the minimizer $T$, we denote the cut (VT,T)(V-T,T) as a min-cut from $s$ to $t$. Given a graph $G$ and $s,t$, a max-flow or minimum $(s,t)$-cut can be computed in $m^{1+o(1)}$ time [chen2022maximum].
2.1 Arborescences
For a directed graph $G$ with source vertex $s$, an arborescence rooted at $s$ is a spanning tree subgraph of $G$ in which every edge is directed away from $s$. Given an integer $k\geq 1$, a $k$-arborescence packing with congestion $\gamma$ is a collection of $k$ arborescences such that each edge appears in at most $\gamma$ arborescences.
2.2 Directed Expanders
Demands.
For a directed graph $G$, a demand $\mathbf{D}:V(G)\times V(G)\to\mathbb{R}_{\geq 0}$ assigns each (ordered) vertex pair a non-negative real number. The demand $\mathbf{D}$ is
  • $\mathscr{C}$-component-constrained for some partition $\mathscr{C}$ of $V(G)$, if for any two vertices $u,v$ belonging to different components $C_{u},C_{v}\in\mathscr{C}$, $\mathbf{D}(u,v)=0$, and
  • $\mathbf{d}$-respecting for some vertex weights $\mathbf{d}:V(G)\to\mathbb{R}_{\geq 0}$, if for each vertex $v$, $\sum_{u}(D(u,v)+D(v,u))\leq\mathbf{d}(v)$.
Routings.
For each pair of vertex $(u,v)$, let ${\cal P}_{u,v}$ denote the collection of simple paths from $u$ to $v$. Let ${\cal P}=\bigcup_{u,v\in V(G)}{\cal P}_{u,v}$ denote the collection of all simple paths. A routing $\mathbf{R}:{\cal P}\to\mathbb{R}_{\geq 0}$ assigns each simple path a non-negative real number. The routing $\mathbf{R}$ has congestion $\gamma$ if for each edge $e\in E(G)$, $\sum_{P\ni e}\mathbf{R}(P)\leq\gamma$. A demand $\mathbf{D}$ can be routed in $G$ with congestion $\gamma$ if there exists a routing $\mathbf{R}$ with congestion $\gamma$ such that for each vertex pair $(u,v)$, $\sum_{P\in{\cal P}_{u,v}}\mathbf{R}(P)=\mathbf{D}(u,v)$.
Weak Flow Expansion.
For a directed graph $G$ with a terminal edge set $E^{\star}\subseteq E(G)$ and a cut edge set $B\subseteq E(G)$, we say $E^{\star}$ is $G\setminus B]-$component-constrained $\phi$-expanding in $G$, if any $(\deg_{E^{\star}})$-respecting $G\setminus B]-$component-constrained demand $\mathbf{D}$ can be routed in $G$ with congestion $1/\phi$.
Lemma 2.1 (Flow Expansion Implies Cut Expansion).
If $E^{\star}$ is $G\setminus B]-$component-constrained $\phi$-expanding in $G$, then for each SCC $C\in[G\setminus B]$ and each cut $(V(G)\setminus T,T)$ such that $\deg_{E^{\star}}(C\cap T)\leq\deg_{E^{\star}}(C)/2$, we have
$\min\{\delta_{G}(T),\rho_{G}(T)\}\geq\phi\cdot\deg_{E^{\star}}(C\cap T).$
2.3 Weighted Graphs
One may also consider weighted graphs $G=(V,E,\mathbf{c})$, with $\mathbf{c}:E\to[W]$, where every edge $e$ is with capacity ${\bf c}(e)$. In weighted graphs, instead of letting $\delta(S)=|E^{+}(S)|$ and $\rho(S)=|E^{-}(S)|$, we let $\delta(S)=\mathbf{c}(E^{+}(S))$ and $\rho(S)=\mathbf{c}(E^{-}(S))$, and the $s$-rooted min-cut is still defined as $\min_{\emptyset\neq S\subseteq V\setminus s}\rho(S)$, which is the minimum weight of a cut.
In Section 3, we use the notion in weighted graphs to conform with [fleischmann2025improved]. In Section 4 and Section 5, we mainly state the results in unweighted graphs. However, all our algorithms will generalize seamlessly to weighted graphs when $W=\mathrm{poly}(n)$.
3 Weak Expander Hierarchy
In this section, we recall tools related to directed expanders. We will exploit directed expander hierarchies to design our approximate min-cut and arborescence packing algorithms.
Definition 3.1 (Expander Hierarchy).
[…]
For intuition, we point out that all SCCs across levels in the hierarchy form a laminar family. Moreover, since we assume that the source vertex $s$ has no incoming edges, $s$ always appears as a singleton SCC in all levels.
Theorem 3.2 ([fleischmann2025improved], Theorem 6.1).
Given a directed, weighted graph $G=(V,E,\mathbf{c})$ with edge capacities $\mathbf{c}$ bounded by $W$, and a terminal edge set $E^{\star}\subseteq E$, there is a randomized algorithm that with high probability finds a set of cut edges $B\subseteq E$ such that
  1. 1.
    $E^{\star}$ is $G\setminus B]-$component-constrained $\phi$-expanding in $G$ for some $\phi=\Omega(1/(\log n\cdot\log^{3}(nW)))$, and
  2. 2.
    the cut size satisfies $\mathbf{c}(B)\leq\mathbf{c}(E^{\star})/2$.
The algorithm runs in $m^{1+o(1)}\log^{3}(nW)$ time.
Remark 3.3.
Theorem 3.2 is not stated identically to Theorem 6.1 in [fleischmann2025improved], but the same argument also yields Theorem 3.2.
Theorem 3.4.
Given a directed, weighted graph $G=(V,E,\textbf{c})$, one can construct a hierarchy $\mathcal{H}$ with $L=O(\log(nW))$ and $\phi=\Omega(1/(\log(n)\log^{3}(nW)))$ in $m^{1+o(1)}\cdot\log^{4}(nW)$ time.
Proof.
Initially, we set $E_{1}=E$. For each level $i=1,2,3,$..., we perform the following steps.
  1. 1.
    Run the directed expander decomposition Theorem 3.2 on $G$ with terminal edge set $E_{i}$, and let $E_{i+1}$ be the returned cut edges.
  2. 2.
    If $E_{i+1}$ is empty, we let $L$ be the current $i$ representing the top level, and then exit the loop. Otherwise, proceed to the next iteration.
The hierarchy $\mathcal{H}$ is exactly $\{E_{i}\mid 1\leq i\leq L\}$ computed above.
Correctness. Consider a level $i$. By Theorem 3.2, $E_{i}$ is $G\setminus E_{i+1}]-$component-constrained $\phi$-expanding in $G$, for $\phi=\Omega(1/(\log(n)\cdot\log^{3}(nW))$. Because $E_{>i}\supseteq E_{i+1}$, we have that $E_{i}$ is $G\setminus E_{>i}]-$component-constrained $\phi$-expanding in $G$ as desired.
The hierarchy has $L=O(\log(nW))$ levels because $\mathbf{c}(E_{i+1})\leq\mathbf{c}(E_{i})/2$ for each $i$, and initially $\mathbf{c}(E_{1})\leq nW$.
Running Time. We run Theorem 3.2 for $L=O(\log(nW))$ levels, so the total running time is $m^{1+o(1)}\log^{4}(nW)$.
4 Approximate rooted min-cut
Given the above hierarchy $\mathcal{H}$ in Definition 3.1, in this section we aim to obtain an $O(L/\phi)$ approximation for the rooted min-cut problem.
See 1.1
We describe our algorithm as follows. First, we consider any level $i$ and any component $C$ (except the singleton $\{s\}$) in $G\setminus E_{>i}]$. Then we consider all edges from $E_{i}$ inside $C$. We sample $O(\log n)$ times: in each iteration we first choose an edge $e$ uniformly at random from $E_{i}[C]$, and then choose a random endpoint $v$ of $e$. For any sampled vertex $v$, we compute the min-cut from $E^{-}(C)$ to $v$, which will correspond to the incoming edges $E^{-}(C_{v})$ for some $C_{v}\subseteq C$. Finally, we output the minimum value among all computed $\rho(C_{v})$. (When $i=0$, every component in $G-E_{>0}]$ is a singleton $C=\{v\}$ and we still find $C_{v}=C$.) A pseudocode is given in Algorithm 1.
Algorithm 1 Approximate $s$-min-cut
[…]
2:for $i=0$ to L, $C\in[G\setminus E_{>i}]$ do
3:  Compute mincut from $E^{-}(C)$ to random $v\leftarrow\text{endpoints of }E_{i}[C]$
4:  Repeat $\Theta(\log n)$ times
5:end for
6:return $\min(\rho(C_{v}))$ for all computed $C_{v}$
Now we prove that the above algorithm will output an $O(L/\phi)$ approximation for the rooted min-cut. Here, we use $\lambda$ to denote the rooted connectivity. From definition, below we suppose $\emptyset\neq T\subseteq V\setminus s$ is a minimizer that $\lambda=\rho(T)$. To prove Theorem 1.1, we aim to find some $C_{v}$ that we compute in the algorithm satisfying $\rho(C_{v})\leq(L/\phi)\lambda=(L/\phi)\rho(T)$.
Start from $T$, our strategy is to construct one candidate set $T_{i}\subseteq T$ for each level $i$ that satisfies $\rho(T_{i})\leq(L/\phi)\rho(T)$ under some conditions. And then, we show that a candidate set $T_{i}$ with $\rho(T_{i})\leq(L/\phi)\rho(T)$ can be captured by some computed $C_{v}$ in the algorithm. We show the formal definition of $\{T_{i}\}$ as follows:
Definition 4.1.
We define $\{T_{i}\}$ inductively from $T_{L}=T$. For each $L>i\geq 0$, we consider an arbitrary topological order between the components in $G-E_{>i}]$. Then we let $C_{i}$ be the first component in the topological order in $G-E_{>i}]$ that intersects $T_{i+1}$, and define $T_{i}=T_{i+1}\cap C_{i}$.
We note that $C_{i}\subseteq C_{i+1}$ for all $i$, since all components in $G-E_{>i}]$ only further partition all components in $G-E_{>{i+1}}]$ and thus $T_{i}=T\cap C_{i}$. Intuitively, $T_{i}$ is the intersection of $T$ and the “first” component $C_{i}$ that intersects $T$. Next, the following theorem gives a condition when $\rho(T_{i})\leq(L/\phi)\rho(T)$ holds.
Theorem 4.2.
For any layer $i$ in $\mathcal{H}$, suppose that for all $k>i$ we have $\mathrm{vol}_{E_{k}}(T_{k})<|E_{k}[C_{k}]|/2$. Then $\rho(T_{i})=O(L\rho(T)/\phi)$.
Proof.
We first claim that, $E^{-}(T_{i})\setminus E^{-}(T)$ contains only edges in $E_{>i}$. Let’s suppose conversely that $E^{-}(T_{i})\setminus E^{-}(T)$ contains some edge $e\in E_{\leq i}$. Since $E^{-}(T_{i})\setminus E^{-}(T)$ only contains edges from $T-T_{i}$ to $T_{i}$, we suppose $e=(u,v)$ with $u\in T-T_{i}$ and $v\in T_{i}$. Now consider the component $C$ in $G-[E_{>i}]$ that intersects $T$ at $u$. Since $C$ has an edge to $C_{i}$, it must be precedent to $C_{i}$ in topological order, which is a contradiction by our definition of $C_{i}$. Thus, we can write as $E^{-}(T_{i})\setminus E^{-}(T)=\cup_{k>i}E_{k}^{-}(T_{i})$.
Next we prove that, for each $k>i$, under the condition that $\mathrm{vol}_{E_{k}}(T_{k})<\mathrm{vol}_{E_{k}}(C_{k})/2$, we have $|E_{k}^{-}(T_{i})|\leq\rho(T)/\phi$. One can observe that, suppose $e\in E_{k}^{-}(T_{i})$, then from the same reasoning as above, both endpoints of $e$ must be in $C_{k}$. Now from Definition 3.1, we know the edges in $E_{k}[C_{k}]$ is $\phi$-expanding. Since $\mathrm{vol}_{E_{k}}(T_{k})<\mathrm{vol}_{E_{k}}(C_{k})/2$, and (VT,T)(V-T,T) is a cut through $E_{k}[C_{k}]$, one can obtain $|E_{k}^{-}(T_{i})|\leq\rho(T)/\phi$ by Lemma 2.1.
Finally, by summing over all $L$ levels, we have $E^{-}(T_{i})\setminus E^{-}(T)\leq L\rho(T)/\phi$ and thus $\rho(T_{i})\leq(L/\phi+1)\rho(T)=O(L\rho(T)/\phi)$. ∎
Now consider the case that $i$ is the largest index that satisfies $\mathrm{vol}_{E_{i}}(T_{i})\geq|E_{i}[C_{i}]|/2$, then $\rho(T_{i})=O(L\rho(T)/\phi)$. Now suppose in the algorithm we have entered the loop with level $i$ and $C=C_{i}$. When $i=0$, $C$ is a singleton and we simply return $\rho(C)=\rho(T_{0})=O(L\rho(T)/\phi)$. Otherwise, we show that given $\mathrm{vol}_{E_{k}}(T_{k})\geq\mathrm{vol}_{E_{k}}(C_{k})/2$, we can capture the set $T_{i}$ by sampling.
Claim 4.3.
When $C=C_{i}$, the algorithm samples some $v\in T_{i}$ with high probability, and $\rho(C_{v})\leq\rho(T_{i})=O(L\rho(T)/\phi)$.
Proof.
By our sampling method, we will sample each vertex $v$ with probability proportional to its degree in $E_{i}$. Now since $\mathrm{vol}_{E_{k}}(T_{k})\geq\mathrm{vol}_{E_{k}}(C_{k})/2$, in every sample, we have at least $1/2$ probability to sample some $v\in T_{i}$. Therefore, the probability that there exists some $v\in T_{i}$ in all $\Theta(\log n)$ samples will be at least $1-(1/2)^{\Theta}(\log n)=1-n^{-\Theta(1)}$. When some $v\in T_{i}$ is sampled, since $(V-T_{i},T_{i})$ is also a valid cut from $E^{-}(C)$ to $v$, we can obtain for the min-cut $C_{v}$ that $\rho(C_{v})\leq\rho(T_{i})$ and thus $\rho(C_{v})=O(L\rho(T)/\phi)$. ∎
So we conclude that the algorithm will output some $\rho(C_{v})=O(L\rho(T)/\phi)$ with high probability, which gives us an $O(L/\phi)$-approximation. Finally, we show the time complexity of our algorithm and prove Theorem 1.1.
Proof of Theorem 1.1.
From above we know the algorithm will output some $\rho(C_{v})=O(L\rho(T)/\phi)$ with high probability. By inserting $L=O(\log n)$ and $\phi=\Omega(1/\log^{4}n)$, one can get $\rho(C_{v})=O(\rho(T)\log^{5}n)$. For time complexity, the hierarchy $\mathcal{H}$ is built in $m^{1+o(1)}$ time in Theorem 3.4. For each layer, all min-cut instances have $\tilde{O}(m)$ edges in total, so the total time of these instances is still $m^{1+o(1)}$. Therefore, the total time for computing $\mathcal{H}$ and resolving $L=O(\log n)$ layers is still $m^{1+o(1)}$. ∎
5 Approximate Arborescence Packing
In this section, we will present our approximate arborescence packing algorithm, formally stated in Theorem 5.2.
5.1 Section Preliminaries
Expander Routing.
In Section 5, we may consider integral demands. An integral demand $\mathbf{D}$ is a multiset of vertex pairs. The notions of being $\mathscr{C}$-component-constrained and $\mathbf{d}$-respecting extend to an integral demand $\mathbf{D}$ by interpreting $\mathbf{D}$ as a function that maps each vertex pair to its multiplicity.
An integral routing $\mathbf{R}$ of an integral demand $\mathbf{D}$ simply maps each $(u,v)\in\mathbf{D}$ into a $u$-$v$ routing path. $\mathbf{R}$ has congestion $\gamma$ if for each edge $e\in E(G)$, the number of routing paths going through $e$ is at most $\gamma$.
Theorem 5.1 (Directed Expander Routing).
Let $G$ be a directed graph with two edge sets $E^{\star},X\subseteq E(G)$ such that $E^{\star}$ is $G\setminus X]-$component-constrained $\phi$-expanding in $G$. Given an integral demand $\mathbf{D}$ on $V(G)$ that is $G\setminus X]-$component-constrained and $(\alpha\cdot\deg_{E^{\star}}^{-})$-respecting for some $\alpha\geq 1$, there is an algorithm that computes an integral routing of $\mathbf{D}$ in $G$ with congestion $\kappa_{\mathrm{route}}\cdot\alpha/\phi$, for some
$\kappa_{\mathrm{route}}=n^{o(1)}.$
The algorithm runs in $(m+|\mathbf{D}|)\cdot n^{o(1)}$ time.
Theorem 5.1 is an extension of almost-linear-time expander routing algorithms in undirected graphs [ghaffari2017distributed, ghaffari2018new, chang2024deterministic]. We will work on a formal proof of this in the full version of the paper.
5.2 The Algorithm
Theorem 5.2.
Let $G=(V,E)$ be a directed graph with a given source vertex $s$ and a parameter $k>0$. Assuming Theorem 5.1, there is an algorithm that computes either
  • a $k$-arborescence packing with congestion $\tilde{O}(\kappa_{\mathrm{route}})=n^{o(1)}$, or
  • a cut $(S,V\setminus S)$ such that $s\in S\subset V(G)$ and $|E^{+}(S)|<k$, certifying that there is no $k$-arborescence packing in $G$.
The running time is $m^{1+o(1)}$.
To obtain a $k$-arborescence packing, we will actually assign each edge $e\in E(G)$ a subset of colors $\Gamma(e)\subseteq[k]$ that satisfies the following.
  1. 1.
    For each color $\gamma\in[k]$, the source $s$ should be able to reach every other vertex using only $\gamma$-colored edges (we say an edge is $\gamma$-colored iff $\gamma\in\Gamma(e)$).
  2. 2.
    Each edge $e\in E(G)$ has at most $\tilde{O}(\kappa_{\mathrm{route}})$ colors, i.e. $|\Gamma(e)|\leq\tilde{O}(\kappa_{\mathrm{route}})$.
Observe that, once we have the above coloring, we can compute $k$ arborescences by for each color $\gamma\in[k]$, picking a DFS tree rooted at $s$ in the subgraph induced by $\gamma$-colored edges. The congestion of these arborescences is clearly bounded by $\tilde{O}(\kappa_{\mathrm{route}})$ because of Property 2.
[…]
$L=O(\log(nW)),\qquad\phi^{-1}=O(\log n\log^{3}(nW)),$
[…]
5.2.1 Invariants at Level $i$
After the computation at level $i$, our algorithm will keep a color assignment $\Gamma_{i}(e)$ for each edge $e\in E(G)$ and $\Gamma_{i}(v)$ for each vertex $v\in V(G)\setminus\{s\}$.
The level-$i$ color assignments should satisfy certain invariants. Before describing the invariants, we start with some notations. For each vertex $v\in V(G)$, we use $C_{i,v}$ to denote the unique SCC in $G\setminus E_{>i}]$ that contains $v$. Moreover, we let $K_{i}(v)$ denote the set of inter-SCC or $E_{>i}$ incoming edges of $v$. That is,
$K_{i}(v):=\{e\in E(G)\mid e=(u,v)\text{\penalty 10000\ s.t.\penalty 10000\ }C_{i,u}\neq C_{i,v}\}\cup E_{>i}^{-}(v),$
and we call the edges in $K_{i}(v)$ the level-$i$ critical incoming edges of $v$. Let
$\Delta_{i}(v)=|K_{i}(v)|$
denote the size of $K_{i}(v)$. Lastly, the following notions of $\gamma$-colored vertices and edges are with respect to the color assignments $\{\Gamma_{i}(e)\}$ and $\{\Gamma_{i}(v)\}$ at level $i$.
The invariants are as follows.
  1. 1.
    For each vertex $v\in V(G)\setminus\{s\}$ and each color $\gamma\in[k]$, $v$ can be reached from a $\gamma$-colored vertex in $C_{i,v}$ using only $\gamma$-colored edges.
  2. 2.
    For each vertex $v\in V(G)\setminus\{s\}$, there is $|\Gamma_{i}(v)|\leq\Delta_{i}(v)\cdot(i+1)$.
  3. 3.
    For each edge $e\in E(G)$, there is $|\Gamma_{i}(e)|\leq 5i^{2}\cdot(\kappa_{\mathrm{route}}/\phi)$.
Intuitively, a vertex $v$ having color $\gamma\in\Gamma_{i}(v)$ means that $v$ acts as a “breakpoint” in the $\gamma$-colored arborescence, and we want to establish reachability from $s$ to $v$ via upper-level routing.
The Base Case.
We start with the bottom level $i=0$, and initialize level-0 color assignments that trivially satisfy the invariants: for each edge $e\in E(G)$, $\Gamma_{0}(e)=\emptyset$; for each vertex $v\in V(G)\setminus\{s\}$, $\Gamma_{0}(v)=[k]$. To see the feasibility of the level-0 color assignments, we first observe each SCC in $G\setminus E_{>0}]$ just contains a single vertex since $E_{>0}=E(G)$. This means $\Delta_{0}(v)=|E^{-}_{G}(v)|$ is just the in-degree of $v$ in $G$.
Invariants 1 and 3 clearly hold. For Invariant 2, if it does not hold, it means there is a vertex $v$ with $\Delta_{0}(v)<k$, which implies a small cut $(V(G)\setminus\{v\},\{v\})$ and we just terminate the algorithm.
The Ending Case: Getting the Final Coloring $\{\Gamma(e)\}$.
Before we describe how to compute the color assignments level by level, we first show how to get the final edge-coloring $\Gamma(e)$ from the top-level color assignments $\{\Gamma_{L}(e)\}$ and $\{\Gamma_{L}(v)\}$. In fact, this is essentially the same as computing an arborescence packing on a DAG.
We start by setting $\Gamma(e)=\Gamma_{L}(e)$ for each edge $e\in E(G)$. For each vertex $v\in V(G)\setminus\{s\}$, observe that $K_{L}(v)$ is exactly the inter-SCC incoming edges of $v$. We distribute the colors $\Gamma_{L}(v)$ over the edges in $K_{L}(v)$ such that each edge receives at most $L+1$ colors. Here, an edge $e\in K_{L}(v)$ receives a subset of colors $Y^{\prime}\subseteq\Gamma_{L}(v)$ means that we add $Y^{\prime}$ into $\Gamma_{i}(e)$. This finishes the computation of the final edge-coloring $\{\Gamma(e)\}$.
Property 1 follows Invariant 1 of the level-$L$ color assignments and the fact that SCCs in $G$ “form” a DAG. Each edge $e$ has
$|\Gamma(e)|\leq|\Gamma_{L}(e)|+L+1\leq O(L^{2}\cdot(\kappa_{\mathrm{route}}/\phi)).$
5.2.2 The Algorithm at Level $i$
Consider a level $1\leq i\leq L$. We now describe the algorithm to compute the level-$i$ color assignments from the level-(i1)(i-1) color assignments. Initially, we set $\Gamma_{i}(v)=\emptyset$ for each $v\in V(G)\setminus\{s\}$ and $\Gamma_{i}(e)=\Gamma_{i-1}(e)$ for each $e\in E(G)$.
Step 1: Allocate “Breakpoint” Colors.
For each SCC $C_{i}\in[G\setminus E_{>i}]$ and each vertex $v\in C_{i}$, recall that $C_{i-1,v}$ is the unique SCC in $G\setminus E_{>i-1}]$ that contains $v$, and clearly $C_{i-1,v}\subseteq C_{i}$. We first partition the level-(i1)(i-1) critical incoming edges of $v$, i.e. $K_{i-1}(v)$, into three sets $E_{X},E_{Y}$ and $E_{Z}$, where
$E_{X}$ $=K_{i}(v),$
$E_{Y}$ $=(E^{-}_{G}(C_{i-1,v})\setminus E^{-}_{G}(C_{i})\setminus E_{>i})\cap E^{-}_{G}(v),$
$E_{Z}$ $=E_{i}^{-}(v)\setminus(E_{X}\cup E_{Y}).$
We make some remarks on this partition. The first set $E_{X}$ is simply the level-$i$ critical incoming edges of $v$, which, by definition, consist of $v$’s incoming edges in $E_{>i}$ and $v$’s incoming edges whose tail-endpoints fall outside $C_{i}$. The second set $E_{Y}$ consists of $v$’s incoming edges whose tail-endpoints fall inside $C_{i}\setminus C_{i-1,v}$, but excludes those inside $E_{>i}$. By definition, $E_{X}$ and $E_{Y}$ are clearly disjoint. By the definition of $K_{i-1}(v)$, the remaining edges $K_{i-1}(v)\setminus(E_{X}\cup E_{Y})$ are all inside $E_{i}^{-}(v)$. Therefore, these three sets $E_{X},E_{Y},E_{Z}$ form a partition of $K_{i-1}(v)$.
Next, we arbitrarily partition $v$’s color $\Gamma_{i-1}(v)$ into three sets $X,Y,Z$, such that
$|X|\leq|E_{X}|\cdot i=\Delta_{i}(v)\cdot i,\qquad|Y|\leq|E_{Y}|\cdot i,\qquad|Z|\leq|E_{Z}|\cdot i.$
  • (Update $\Gamma_{i}(v)$) For the colors in $X$, we add them to $\Gamma_{i}(v)$.
  • (Update $\Gamma_{i}$(e)) For the colors in $Y$, we distribute them over edges in $Y$ so that each edge in $Y$ receives at most $i$ colors from $Y$.
The remaining colors in $Z$ will be handled in the following steps, and we use $Z_{v}$ refer to the $Z$ for this vertex $z$.
Step 2: Max Flows.
For each SCC $C_{i}\in[G\setminus E_{>i}]$, we will run a single-commodity max-flow instance in $G[C_{i}]$. The instance has source function $\Delta_{C_{i}}:=(\Delta_{i})_{\mid C_{i}}$, i.e., the restriction of $\Delta_{i}$ on vertices $C_{i}$, and sink function $\nabla_{C_{i}}:=(i\cdot\deg^{-}_{E_{i}})_{\mid C_{i}}$, i.e., each vertex $v\in C_{i}$ has $\nabla_{C_{i}}(v)$ equal to $i$ times the number of $v$’s incoming edges in $E_{i}$. Note that, if we let
$Z_{C_{i}}:=\bigcup_{v\in C_{i}}Z_{v}$
be the union of the $Z$-colors of all $v\in C_{i}$, then the total sink $|\nabla_{C_{i}}|:=\sum_{v\in C_{i}}\nabla_{C_{i}}(v)$ satisfies
$|\nabla_{C_{i}}|=i\cdot\sum_{v\in C_{i}}|E^{-}_{i}(v)|\geq\sum_{v\in C_{i}}|Z_{v}|\geq|Z_{C_{i}}|$
by the definition of $Z_{v}$
Lemma 5.3.
By calling an exact max-flow algorithm on $G[C_{i}]$ once, we can obtain either
  • (Cut Case) a non-empty set $C^{\star}\subseteq C_{i}$ such that $|E^{-}(C^{\star})|<k$, or
  • (Flow Case) a collection of $|Z_{C_{i}}|$ edge-disjoint paths in $G[C_{i}]$, denoted by ${\cal P}_{C_{i}}$, such that: for each vertex $u\in C_{i}$, at most $\Delta_{C_{i}}(u)=\Delta_{i}(u)$ paths start at $u$; and for each vertex $v\in C_{i}$, at most $\nabla_{C_{i}}(v)$ paths end at $v$.
We are not going to prove Lemma 5.3 formally since it is standard. Roughly speaking, if the max-flow value is less than $\min\{k,|\nabla_{C_{i}}|\}$, then it falls into the cut case. Otherwise, it falls into the flow case by decomposing the max-flow of value at least $\min\{k,|\nabla_{C_{i}}|\}\geq|Z_{C_{i}}|$.
By Lemma 5.3, if we obtain a set $C^{\star}$, then $(V(G)\setminus C^{\star},C^{\star})$ is a desired small cut, and thus we can terminate the whole algorithm. Otherwise, we obtain the edge-disjoint paths ${\cal P}_{C_{i}}$. We assign each of them to each color in $Z_{C_{i}}$.
  • (Update $\Gamma_{i}(e)$) For each color $\gamma\in Z_{C_{i}}$, we assign the color $\gamma$ to each edge $e$ on the corresponding path $P_{C_{i},\gamma}$ (i.e. add $\gamma$ to $\Gamma_{i}(e)$ for each $e\in P_{C_{i},\gamma}$)
  • (Update $\Gamma_{i}(v)$) For the starting vertex of $P_{C_{i},\gamma}$, say $u$, we add the color $\gamma$ to $\Gamma_{i}(u)$.
Lastly, we refer to the ending vertex of $P_{C_{i},\gamma}$ as the $\gamma$-colored leader of $C_{i}$, denoted by $w_{C_{i},\gamma}$.
Step 3: Expander Routing.
For each SCC $C_{i}\in[G\setminus E_{>i}]$ and each color $\gamma\in Z_{C_{i}}$, we refer to a vertex $v\in C_{i}$ with $\gamma\in Z_{v}$ as a $\gamma$-colored breakpoint, and we let
$Z^{-1}_{C_{i}}(\gamma)=\{v\in C_{i}\mid\gamma\in Z_{v}\}$
collect all $\gamma$-colored breakpoints of $C_{i}$. Then, we fix an arbitrary chain demand $\mathbf{D}_{C_{i},\gamma}$ over the $\gamma$-colored leader $w_{C_{i},\gamma}$ and breakpoints $Z^{-1}_{C_{i}}(\gamma)$ of $C_{i}$. That is, pick an arbitrary order of vertices in $Z^{-1}_{C_{i}}(\gamma)$, say $v_{1},v_{2},...,v_{q}$ (where $q=|Z^{-1}_{C_{i}}(\gamma)|$), and then create $q$ demand pairs $(w_{C_{i},\gamma},v_{1})$, $(v_{1},v_{2})$, …, $(v_{q-1},v_{q})$.
Let $\mathbf{D}_{i}=\sum_{C_{i},\gamma}\mathbf{D}_{C_{i},\gamma}$ be the total demand set. The following Lemma 5.4 enables us to route $\mathbf{D}_{i}$ via expander routing.
Lemma 5.4.
The total demand $\mathbf{D}_{i}$ is $(3i\cdot\deg^{-}_{E_{i}})$-respecting and $G\setminus E_{>i}]-$component-constrained.
Proof.
First, $\mathbf{D}_{i}$ is clearly $G\setminus E_{>i}]-$component-constrained by definition. To see that $\mathbf{D}_{i}$ is $(3i\cdot\deg^{-}_{E_{i}})$-respecting, we will show that each vertex $v\in V(G)$ appears in at most $3i\cdot|E^{-}(v)|$ demand pairs. First, $v$ can serve as the $\gamma$-colored leaders of $C_{i,v}$ for at most $\nabla_{C_{i}}(v)=i\cdot|E^{-}(v)|$ colors by Lemma 5.3. This contributes at most $i\cdot|E^{-}_{i}(v)|$ appearances. Second, $v$ is a $\gamma$-colored breakpoint of $C_{i,v}$ for each color $\gamma\in Z_{v}$ (note that $|Z_{v}|\leq i\cdot|E^{-}_{i}(v)|$), and each breakpoint contributes two appearances. ∎
By Lemma 5.4 and Theorem 5.1, we can obtain a routing $\mathbf{R}_{i}$ of $\mathbf{D}_{i}$ in $G$ with
$\text{congestion\penalty 10000\ }(3i)\cdot\kappa_{\mathrm{route}}/\phi.$
Therefore, for each chain demand $\mathbf{D}_{C_{i},\gamma}$ and each demand pair $(w_{1},w_{2})\in\mathbf{D}_{C_{i},\gamma}$, we have an associated routing path $R$ in $G$ from $w_{1}$ to $w_{2}$.
  • (Update $\Gamma_{i}(e)$) For each edge $e$ on $R$, we add color $\gamma$ to $\Gamma_{i}(e)$.
This finishes the computation of $\{\Gamma_{i}(e)\mid e\in E(G)\}$ and $\{\Gamma_{i}(v)\mid v\in V(G)\setminus\{s\}\}$.
Proof of the Invariants.
We now show that the level-$i$ color assignments satisfy all the invariants.
Invariant 1. First of all, observe that SCCs $C_{i-1}\in[G\setminus E_{>i-1}]$ admit a topological order $\pi_{i-1}$, such that for each edge $e=(u,v)\in E(G)\setminus E_{>i-1}$ connecting two different SCCs $C_{i-1,u}$ and $C_{i-1,v}$, the SCC $C_{i-1,u}$ precedes $C_{i-1,v}$ on $\pi_{i-1}$. Moreover, the SCC $C_{i-1,s}$ containing the source vertex $s$ is the first in $\pi_{i-1}$.
Now, consider a vertex $v\in V(G)\setminus\{s\}$ and a color $\gamma\in[k]$. To avoid clutter, let $C_{i}$ be the SCC $C_{i,v}\in[G\setminus E_{>i}]$ containing $v$. We want to show that $v$ can be reached from a $\gamma$-colored (w.r.t. $\Gamma_{i}(\cdot)$) vertex in $C_{i}$, using only $\gamma$-colored edges. By the invariants of the level-(i1)(i-1) color assignments, we know that $v$ can be reached by a vertex $u\in C_{i-1,v}$ s.t. $\gamma\in\Gamma_{i-1}(u)$ using only $\gamma$-colored edges. Recall that in Step 1, we partition $\Gamma_{i-1}(u)$ into three sets $X_{u},Y_{u},Z_{u}$.
  • Suppose $\gamma\in X_{u}$. Then according to the update of $\Gamma_{i}(u)$ in Step 1, $\gamma$ is added to $\Gamma_{i}(u)$, and we are done.
  • Suppose $\gamma\in Z_{u}$. Following the routing paths of the chain demand $\mathbf{D}_{C_{i},\gamma}$, $u$ can be reached from the $\gamma$-colored leader $w_{C_{i},\gamma}$ of $C_{i}$ using only $\gamma$-colored edges. Moreover, following the flow path $P_{C_{i},\gamma}$, $w_{C_{i},\gamma}$ can be reached from the starting vertex of $P_{C_{i},\gamma}$ (which is a $\gamma$-colored vertex) using only $\gamma$-colored edges.
  • Suppose $\gamma\in Y_{u}$. In Step 1, $\gamma$ is assigned to an incoming edge $(v^{\prime},u)$ of $u$ such that $v^{\prime}\in C_{i}\setminus C_{i-1,v^{\prime}}$. This time, we repeat the above argument for the vertex $v^{\prime}$ and the color $\gamma$. Note that we make progress each time, as $C_{i-1,v^{\prime}}$ precedes $C_{i-1,v}$ in the order $\pi_{i}$, so finally we will reach the above two cases which find a $\gamma$-colored vertex in $C_{i}$ that can reach $v$ using only $\gamma$-colored edges.
Invariant 2. Consider a vertex $v\in V(G)\setminus\{s\}$. Initially, $\Gamma_{i}(v)$ is empty. In Step 1, the size of $\Gamma_{i}(v)$ increases by at most $\Delta_{i}(v)\cdot i$. In Step 2, the size of $\Gamma_{i}(v)$ increases by at most $\Delta_{i}(v)$ because of Lemma 5.3. Hence $|\Gamma_{i}(v)|\leq\Delta_{i}(v)\cdot(i+1)$ at the end.
Invariant 3. Consider an edge $e\in E(G)$. Initially, $|\Gamma_{i}(e)|=|\Gamma_{i-1}(e)|\leq 5(i-1)^{2}\cdot\kappa_{\mathrm{route}}/\phi$. In Steps 1, 2 and 3, $|\Gamma_{i}(e)|$ increases by at most $i$, 11 and $(3i)\cdot\kappa_{\mathrm{route}}/\phi$ respectively. Therefore,
$|\Gamma_{i}(e)|\leq 5(i-1)^{2}\cdot\kappa_{\mathrm{route}}/\phi+i+1+(3i)\cdot\kappa_{\mathrm{route}}/\phi\leq 5i^{2}\cdot\kappa_{\mathrm{route}}/\phi.$
5.2.3 Running Time Analysis
The bottlenecks are the expander hierarchy algorithm (Theorem 3.4), exact max-flow algorithm (Lemma 5.3), and the expander routing (Theorem 5.1). All of them incur $n^{o(1)}$ overheads, which dominate other polylogarithmic overheads. Hence the total running time is $m^{1+o(1)}$.